OpenStudy (anonymous):

1/(x^2-7x+10) = x/(x-5) + 1(x-2)

OpenStudy (campbell_st):

whats the question..?

OpenStudy (anonymous):

Solve for x

OpenStudy (campbell_st):

ok... the left hand side in factorised form is \[\frac{1}{(x -2)(x -5)} \] the write hand side needs a common denominator \[(x -2)(x-5)\] you can rewrite the right hand side as \[\frac{x}{(x -5)} + \frac{1}{(x -2)} = \frac{x (x -2) + (x -5)}{(x-2)(x-5)}\] since both sides now have the same denominator then equate he numerators 1 = x(x-2) + x - 5 simplifying \[1 = x^2 - 2x + x - 5 \] or \[x^2 - x - 6 = 0\] you can solve this equation by factorising hope this helps

OpenStudy (anonymous):

Oh , lol well thank you .. it did !

OpenStudy (campbell_st):

well done